f you deposit $2294 in your bank account today and wait 18 years, how much money will be in your account if the interest rate is 2.7% compounded monthly?
1
Expert's answer
2020-10-23T11:48:04-0400
"2294\times(1+0.027)^{18 \times 12}=724117.812"